# FedSDAF: Leveraging Source Domain Awareness for Enhanced Federated Domain Generalization

### v3

@@ Abstract
Traditional Federated Domain Generalization (FedDG) methods focus on learning domain-invariant features or adapting to unseen target domains, often overlooking the unique knowledge embedded within the source domain, especially in strictly isolated federated learning environments. Through experimentation, we discovered a counterintuitive phenomenon.: features learned from a complete source domain have superior generalization capabilities compared to those learned directly from the target domain. This insight leads us to propose the Federated Source Domain Awareness Framework (FedSDAF), the first systematic approach to enhance FedDG by leveraging source domain-aware features. FedSDAF employs a dual-adapter architecture that decouples "local expertise" from "global generalization consensus". A Domain-Aware Adapter, retained locally, extracts and protects the unique discriminative knowledge of each source domain, while a Domain-Invariant Adapter, shared across clients, builds a robust global consensus. To enable knowledge exchange, we introduce a Bidirectional Knowledge Distillation mechanism that facilitates efficient dialogue between the adapters. Extensive experiments on four benchmark datasets (OfficeHome, PACS, VLCS, DomainNet) show that FedSDAF significantly outperforms existing FedDG methods.The source code is available at https://github.com/pizzareapers/FedSDAF.
1 Introduction
@@
The synergy between federated learning (FL) and large-scale pre-trained models [2, 3] has ushered in a new era of distributed learning paradigms for privacy-sensitive applications. While foundational architectures, particularly Transformers, demonstrate remarkable cross-task adaptability [1, 4] through parameter efficient techniques, conventional FL frameworks [5, 6] encounter a fundamental limitation: their underlying independent and identically distributed (i.i.d.) assumption stands in stark contrast to the inherent multi-source domain heterogeneity in real-world scenarios [7, 8, 9]. This misalignment presents a dual challenge: not only do client-specific data emerge from distinct domain distributions, but models must also simultaneously address inter-client domain shifts while maintaining robust generalization capabilities for unseen target domains during deployment. These intertwined challenges constitute the core problem in Federated Domain Generalization (FedDG).
@@
Standard domain generalization (DG) methods have advanced through Domain-invariant representation learning [14, 15] and Feature disentanglement [16]. Both Feng et al. [12] and Li et al. [13] have adopted the Feature Disentanglement approach, which separates features into domain-invariant and domain-specific components, utilizing domain-aware features from the target domain to enhance the model’s generalization ability in text classification and target detection tasks. However, these studies primarily focus on leveraging domain-aware features in the target domain, without fully exploring the potential value of domain-aware features in the source domain.
@@
2 Related Work
2.1 Federated Learning and Domain Generalization
@@
2.2 Federated Domain Generalization
@@
3 Methodology
3.1 Problem Setup
In the Federated Domain Generalization (FedDG) setting, we consider a scenario involving $K$ independent clients. Each client, $C_{k}$, exclusively holds a local source dataset $D_{k}=\{(x_{ik},y_{ik})\}_{i=1}^{n_{k}}$ that is drawn from a unique probability distribution $P_{k}(X,Y)$, such that $P_{k}\neq P_{j}$ for any $k\neq j$. The foundational constraint of this setting is: Data Isolation, raw data sharing among clients is strictly prohibited due to privacy concerns. Clients can only collaborate to train a model by exchanging non-sensitive information, such as model parameters or updates, via a central server. The primary objective in FedDG is to leverage the diverse, isolated source domains $\{D_{k}\}_{k=1}^{K}$ to learn a single, robust global model $f_{\theta_{g}}$ that performs well on an entirely unseen target domain $D_{t}$. This target domain is drawn from a different distribution $P_{t}$, where $P_{t}\neq P_{k}$ for all source clients $k\in\{1,\ldots,K\}$. The global model $f_{\theta_{g}}$ is produced by aggregating the local models trained on each client’s private data. This constrained optimization problem is formally expressed as minimizing the expected risk on the target domain:
@@ 3.1 Problem Setup
3.2 Motivational Study: The Case for Source-Domain Awareness
@@
Table 1: Performance (%) of adapters trained on a single source domain (rows) and evaluated on a single target domain (columns).
@@
Results and Principled Interpretation. The results, presented in Table 1, reveal a clear and theoretically grounded trend. In almost all cases, the Source-Aware adapter (off-diagonal values) outperforms the target-aware adapter (diagonal values) when evaluated on the target domain. For instance, when targeting the Sketch domain, the adapter trained on Art Painting (97.07%) significantly surpasses the one trained on Sketch itself (94.94%).
@@ 3.3 Proposed Method
3.3.1 Knowledge Integrator Adapter (KIA)
@@
(1) Domain-Invariant Adapter: The Domain-Invariant Adapter $A_{di}$ aims to capture stable and generalizable representations across all source domains. At each federated communication round $r$, the server maintains a global domain-invariant adapter $A_{di}^{r}$, which is distributed to all clients for local fine-tuning. Each client $k$ optimizes a local copy of this adapter, denoted as $\hat{A}_{di}^{k,r}$, on its private dataset. After local training, the server aggregates these adapters into an updated global adapter using a weighted averaging scheme, which is essential for maintaining a robust representation that can effectively generalize across various domains.
(2) Domain-Aware Adapter: Before the first round of communication, each client initializes a domain-aware adapter $A_{dw}$ locally. Unlike the domain-invariant adapter, this domain-aware adapter is updated strictly locally throughout the training process to force it to capture the domain-aware features of each domain.. Its role is to capture subtle, client-specific knowledge from the local domain, thus preserving rich domain-specific information. The domain-aware adapter is crucial for enhancing the model’s ability to adapt to the specific characteristics of the client’s data.
The Knowledge Integrator Adapter (KIA), which integrates both the Domain-Invariant Adapter and the Domain-Aware Adapter (as illustrated in Fig. 1), is inserted following the Feedforward Neural Networks (FFNs) layer of the neural network. To effectively integrate these two sets of features, we employ a Multi-Head Self-Attention (MHSA) mechanism. Let $h$ denote the intermediate representation output from the backbone model. The integrated feature representation on client $k$ is computed as follows:
$\hat{A}_{dw}\left(h\right)=MHSA\left(LN\left[A_{dw}\left(h\right);h\right]\right)$ (1)
@@
This integration ensures that in each round of communication, the model effectively assimilates knowledge from the source domain while leveraging the rich contextual information inherent in the domain-aware features. By utilizing MHSA, the model can dynamically weigh the importance of various features from both the source domain and the base model’s outputs, facilitating the formation of more adaptive domain-aware features.
@@
3.3.2 Bidirectional Knowledge Distillation (BKD)
The bidirectional knowledge distillation process serves as the core mechanism within the FedSDAF framework, designed to facilitate effective knowledge sharing among various clients. This process enables the model to transfer and integrate knowledge across different source domains, thereby achieving a comprehensive utilization of knowledge from multiple sources. Specifically, the bidirectional knowledge distillation involves the following steps: First, two Kullback-Leibler (KL) divergences are computed using the following formulas:
$L_{di}^{KL}=KL(z_{di}\circ(x)\parallel\hat{z}_{dw}\circ(x))$ (3)
$L_{dw}^{KL}=KL(\hat{z}_{dw}\circ(x)\parallel z_{di}\circ(x))$ (4)
Here, KL denotes the Kullback-Leibler divergence, while $z_{di}$ and $\hat{z}_{dw}$ represent the prediction logits of the base model injected with $A_{di}$ and $\hat{A}_{dw}$, respectively. Through this setup, bidirectional knowledge distillation allows clients to exchange knowledge, facilitating learning and adaptation across different source domains. By sharing prediction logits, the model is able to acquire additional domain information from the server pertaining to other domains. This interaction of information is crucial in compensating for knowledge gaps arising from data isolation.
During the actual training process, we employ a distillation loss in conjunction with Cross-Entropy Loss (CEL), introducing a weighting factor to maintain a balance between the two losses, Figure 2 illustrates the process of bi-directional distillation:
@@ 3.3.2 Bidirectional Knowledge Distillation (BKD)
3.3.3 Optimization process
The optimization process unfolds across the server and clients. On the server, the global domain-invariant adapter $A_{di}$ is initialized. In each communication round $r$, the server distributes the current global adapter $A_{di}^{r}$ to a subset of clients. Each participating client $k$ receives $A_{di}^{r}$ and performs local training on its dataset $\mathcal{D}_{k}$. During local training, the client updates its local copy of the domain-invariant adapter, $\hat{A}_{di}^{k,r}$, and its private domain-aware adapter, $A_{dw}^{k}$, using the bidirectional knowledge distillation loss described in Equation (5). After local training, each client sends its updated adapter $\hat{A}_{di}^{k,r}$ back to the server. The server then aggregates the received adapters to produce the next round’s global adapter, $A_{di}^{r+1}$, using standard federated averaging:
@@ 3.3.3 Optimization process
where $n_{k}$ is the number of samples on client $k$ and $N=\sum_{k}n_{k}$ is the total number of samples across all clients. This iterative process allows the framework to build a robust global consensus ($A_{di}$) while leveraging the rich, localized knowledge preserved within each client’s $A_{dw}$.
4 Experimental Results
4.1 Datasets and Baselines
  1. 1.
  2. 2.
  3. 3.
Table 2: Performance Comparison on PACS, VLCS, and OfficeHome Datasets (%)
Methods PACS VLCS OfficeHome
A C P S Avg. C L V S Avg. A C P R Avg.
Centralized learning based Domain Generalization Methods
CNN-based FedDG Methods (backbone: Resnet-50)
PEFT Methods (backbone: ViT-B/16)
Table 3: Performance Comparison on DomainNet Datasets (%)
Methods DomainNet
C I P Q R S Avg.
FedCLIP $74.12$ $48.36$ $68.49$ $31.73$ $80.52$ $58.62$ $60.31$
PromptFL $76.53$ $51.72$ $70.86$ $34.21$ $81.68$ $68.37$ $63.90$
FedAPT $77.02$ $51.45$ $70.36$ $49.62$ $86.64$ $68.43$ $67.25$
FedPR $75.49$ $51.96$ $71.42$ $35.98$ $82.67$ $69.43$ $64.49$
FedMaPLe $78.61$ $65.23$ $71.89$ $43.46$ $86.32$ $72.46$ $69.67$
PLAN $79.51$ 66.42 $72.11$ $48.83$ 86.72 $72.69$ $71.05$
FedSDAF 82.59 $62.95$ 85.12 54.16 $84.99$ 74.14 73.99
4.2 Implementation Details
In our implementation of FedSDAF, we employed the ViT-B16 architecture as the backbone network. The federated learning process was configured to span 200 communication rounds, with each client performing 30 local epochs per round. To accommodate the diverse characteristics of the benchmark datasets, we utilized dataset-specific batch sizes: 128 for both PACS and Office Home datasets, 64 for VLCS, and 1024 for the more extensive DomainNet. For optimization, we adopted the Scaffold optimizer with an initial learning rate of 0.001 to facilitate effective coordination of client model updates. To regulate the learning rate throughout the training process, we implemented PyTorch’s StepLR scheduler with a step size of 0.2 and a decay factor of 0.1, ensuring appropriate learning rate adjustments during model convergence.
4.3 Main Results
Tables 2 to 3 show the comparison of our proposed FedSDAF method with other methods on PACS, VLCS, Officehome and DomainNet, respectively.
In our experimental evaluations, we achieved state-of-the-art (SOTA) results across all four mainstream datasets: PACS, VLCS, OfficeHome, and DomainNet. This accomplishment underscores the robust generalization capabilities of our proposed Federated Source Domain Awareness Framework (FedSDAF). Notably, our method demonstrated significant advancements on the S-domain of the PACS dataset, where previous approaches had struggled to deliver satisfactory outcomes. Specifically, our framework achieved an impressive accuracy of 98.62%, reflecting a notable improvement of 3.6% over prior methods.
4.4 Ablation Experiment
Table 4: Ablation Study for Different Components
BKD MHSA $\hat{A}_{di}$ $\hat{A}_{dw}$ PACS OfficeHome VLCS
$\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 98.52 87.63 91.38
$\times$ $\checkmark$ $\checkmark$ $\checkmark$ 97.46 85.81 86.19
$\checkmark$ $\times$ $\checkmark$ $\checkmark$ 97.13 85.42 87.20
$\checkmark$ $\times$ $\checkmark$ $\times$ 95.47 83.60 83.98
$\times$ $\times$ $\checkmark$ $\times$ 95.46 82.17 83.51
$\times$ $\times$ $\times$ $\times$ 93.29 78.70 82.69
To illustrate the significance of the various components utilized in the Federated Source Domain Awareness Framework (FedSDAF), we conducted an ablation study across three benchmark datasets. The results are presented in Table 4.
Optimal model performance was attained when all components were active, demonstrating the synergistic integration of features that collectively enhance the model’s generalization capacity. In contrast, disabling both the Bidirectional knowledge distillation (BKD) and the Multihead Self-Attention mechanism (MHSA) led to a significant decline in performance, highlighting the critical roles these components play in facilitating information exchange and feature extraction. The observed drop in accuracy indicates that, without these mechanisms, the model struggles to maintain the understanding necessary for effective generalization. This finding underscores the necessity of collaborative learning strategies, such as BKD, in optimizing model performance within federated contexts characterized by data isolation.
4.5 Convergence Analysis
To validate the convergence of our FedSDAF framework, we compared it against the FedAVG (Adapter) baseline on the PACS dataset. As shown in Figure 3, FedSDAF converges significantly faster and more stably. Within the first 10 rounds, FedSDAF already achieves a 5.8-7.2% accuracy advantage. This rapid start is driven by the local domain-aware adapter’s MHSA mechanism, which accelerates the extraction of domain-specific features, whereas FedAVG’s domain-agnostic aggregation lags behind. By round 40, FedSDAF’s lead widens to a stable 9.4-13.8% accuracy margin. This superior stability is a direct result of our bidirectional knowledge distillation, which systematically fuses complementary knowledge across domains. This dynamic process overcomes the limitations of static aggregation strategies that are typically hindered by domain heterogeneity, ensuring a consistently faster and more robust convergence.
4.6 Hyperparameter Analysis
Figure 4: Effects of $\alpha$.
We evaluate the robustness of FedSDAF through systematic analysis of the loss balancing coefficient $\alpha$ (Eq.7) across the entire PACS benchmark. FedSDAF consistently achieves accuracy exceeding 97% when $\alpha$ ranges from 0.3 to 5, with minimal variance (±0.8%). Notably, even with $\alpha$=0.1 (representing reduced distillation influence), the model maintains 96.2% accuracy, demonstrating the inherent stability of our dual-adapter architecture. Optimal performance is observed at $\alpha$=1, where domain-specific knowledge preservation and cross-client fusion reach an equilibrium. When $\alpha$ exceeds 5, a slight performance decline occurs (97.1% at $\alpha$=10), as the over-strengthened global distillation signal marginally inhibits local adaptation capabilities. These experimental results confirm FedSDAF’s robustness to hyperparameter selection, requiring only the straightforward choice of $\alpha$=1 to achieve peak performance without extensive tuning.
4.7 Communication Cost Analysis
The communication cost per round is dominated by two components: (1) the upload/download of the lightweight Domain-Invariant Adapter ($\hat{A}_{di}$), and (2) the exchange of logits for the Bidirectional Knowledge Distillation (BKD) process. As shown in Figure 5, the size of the logits is negligible (0.003 MB). The total estimated communication cost for FedSDAF per client per round is approximately 2.270 MB. This is not only comparable to other PEFT-based methods like FedClip (2.010 MB) but is also an order of magnitude smaller than full-model aggregation methods like FedAvg (13.650 MB). This analysis demonstrates that FedSDAF remains highly communication-efficient, highlighting its practical potential for deployment in real-world federated learning scenarios.
5 Conclusion
Appendix A Algorithm Details
As depicted in Algorithm 1, on the client side, each client initializes its local copy $\hat{A}_{di}^{k,r}$ corresponding to the received global domain-invariant adapter $A_{di}^{r}$. The client subsequently optimizes its local domain-aware adapter $A_{dw}^{k,r}$ and the local replica $\hat{A}_{di}^{k,r}$ through the bidirectional knowledge distillation mechanism. During this optimization, the client computes its training error $\hat{E}_{D_{k}}^{k,r}(A_{di}^{k,r})$ and, after each round, uploads the generalization gap $G_{k}^{r}$ (if $r>0$) along with the updated domain-invariant adapters $A_{di}^{k,r}$ back to the server. This systematic updating process enables the FedSDAF method to effectively leverage source domain knowledge, thereby enhancing the model’s generalization performance in unseen target domains.
@@ Appendix A Algorithm Details
1:$K$: Number of clients; $\{D_{k}\}_{k=1}^{K}$: Client datasets; $R$: Communication rounds
[…]
@@
Appendix B Supplementary Experiment
B.1 Fine-grained Ablation Study on Hard Domains

### v4

@@ Abstract
Traditional Federated Domain Generalization (FedDG) methods focus on learning domain-invariant features or adapting to unseen target domains, often overlooking the unique knowledge embedded within the source domain, especially in strictly isolated federated learning environments. Through experimentation, we discovered a counterintuitive phenomenon: features learned from a complete source domain have superior generalization capabilities compared to those learned directly from the target domain. This insight leads us to propose the Federated Source Domain Awareness Framework (FedSDAF), the first systematic approach to enhance FedDG by leveraging source domain-aware features. FedSDAF employs a dual-adapter architecture that decouples "local expertise" from "global generalization consensus." A Domain-Aware Adapter, retained locally, extracts and protects the unique discriminative knowledge of each source domain, while a Domain-Invariant Adapter, shared across clients, builds a robust global consensus. To enable knowledge exchange, we introduce a Bidirectional Knowledge Distillation mechanism that facilitates efficient dialogue between the adapters. Extensive experiments on four benchmark datasets (OfficeHome, PACS, VLCS, and DomainNet) show that FedSDAF significantly outperforms existing FedDG methods. The source code is available at https://github.com/pizzareapers/FedSDAF.
I Introduction
Figure 1: Illustration of domain generalization.
@@ I Introduction
The synergy between federated learning (FL) and large-scale pre-trained models [2, 3] has ushered in a new era of distributed learning paradigms for privacy-sensitive applications. While foundational architectures, particularly Transformers, demonstrate remarkable cross-task adaptability [1, 4] through parameter-efficient techniques, conventional FL frameworks [5, 6] encounter a fundamental limitation: their underlying independent and identically distributed (i.i.d.) assumption stands in stark contrast to the inherent multi-source domain heterogeneity in real-world scenarios [7, 8, 9]. This misalignment presents a dual challenge: not only do client-specific data emerge from distinct domain distributions, but models must also simultaneously address inter-client domain shifts while maintaining robust generalization capabilities for unseen target domains during deployment. These intertwined challenges constitute the core problem in Federated Domain Generalization (FedDG).
@@
Standard domain generalization (DG) methods have advanced through domain-invariant representation learning [14, 15] and feature disentanglement [16]. Both [12] and [13] have adopted the feature disentanglement approach, which separates features into domain-invariant and domain-specific components, utilizing domain-aware features from the target domain to enhance the model’s generalization ability in text classification and target detection tasks. However, these studies primarily focus on leveraging domain-aware features in the target domain, without fully exploring the potential value of domain-aware features in the source domain.
@@
II Related Work
II-A Federated Learning and Domain Generalization
@@
II-B Federated Domain Generalization
@@
III Methodology
III-A Problem Setup
In the Federated Domain Generalization (FedDG) setting, we consider a scenario involving $K$ independent clients. Each client, $C_{k}$, exclusively holds a local source dataset $D_{k}=\{(x_{ik},y_{ik})\}_{i=1}^{n_{k}}$ that is drawn from a unique probability distribution $P_{k}(X,Y)$, such that $P_{k}\neq P_{j}$ for any $k\neq j$. The foundational constraint of this setting is Data Isolation: raw data sharing among clients is strictly prohibited due to privacy concerns. Clients can only collaborate to train a model by exchanging non-sensitive information, such as model parameters or updates, via a central server. The primary objective in FedDG is to leverage the diverse, isolated source domains $\{D_{k}\}_{k=1}^{K}$ to learn a single, robust global model $f_{\theta_{g}}$ that performs well on an entirely unseen target domain $D_{t}$. This target domain is drawn from a different distribution $P_{t}$, where $P_{t}\neq P_{k}$ for all source clients $k\in\{1,\ldots,K\}$. The global model $f_{\theta_{g}}$ is produced by aggregating the local models trained on each client’s private data. This constrained optimization problem is formally expressed as minimizing the expected risk on the target domain:
@@ III-A Problem Setup
III-B Motivational Study: The Case for Source-Domain Awareness
@@
TABLE I: Performance (%) of adapters trained on a single source domain (rows) and evaluated on a single target domain (columns).
@@
Results and Principled Interpretation. The results, presented in Table I, reveal a clear and theoretically grounded trend. In almost all cases, the source-aware adapter (off-diagonal values) outperforms the target-aware adapter (diagonal values) when evaluated on the target domain. For instance, when targeting the Sketch domain, the adapter trained on Art Painting (97.07%) significantly surpasses the one trained on Sketch itself (94.94%).
@@ III-C Proposed Method
III-C1 Knowledge Integrator Adapter (KIA)
@@
Domain-Invariant Adapter $A_{di}$: The $A_{di}$ aims to capture stable and generalizable representations across all source domains. At each federated communication round $r$, the server maintains a global domain-invariant adapter $A_{di}^{r}$, which is distributed to all clients for local fine-tuning. Each client $k$ optimizes a local copy of this adapter, denoted as $\hat{A}_{di}^{k,r}$, on its private dataset. After local training, the server aggregates these adapters into an updated global adapter using a weighted averaging scheme, which is essential for maintaining a robust representation that can effectively generalize across various domains.
Domain-Aware Adapter $A_{dw}$: Before the first round of communication, each client initializes a $A_{dw}$, locally. Unlike the domain-invariant adapter, this adapter is updated strictly locally throughout the training process to force it to capture the domain-aware features of each domain. Its role is to capture subtle, client-specific knowledge from the local domain, thus preserving rich domain-specific information. The domain-aware adapter is crucial for enhancing the model’s ability to adapt to the specific characteristics of the client’s data.
The KIA, which integrates both $A_{di}$ and $A_{dw}$ (as illustrated in Fig. 2), is inserted following the Feedforward Neural Networks (FFNs) layer of the neural network. To effectively integrate these two sets of features, we employ a Multi-Head Self-Attention (MHSA) mechanism. Let $h$ denote the intermediate representation output from the backbone model. The integrated feature representation on client $k$ is computed as follows:
$\hat{A}_{dw}(h)=\text{MHSA}(\text{LN}[A_{dw}(h);h])$ (1)
@@
This integration $h^{\prime}$ ensures that in each round of communication, the model effectively assimilates knowledge from the source domain while leveraging the rich contextual information inherent in the domain-aware features. By utilizing MHSA, the model can dynamically weigh the importance of various features from both the source domain and the base model’s outputs, facilitating the formation of more adaptive domain-aware features.
@@
III-C2 Bidirectional Knowledge Distillation (BKD)
The bidirectional knowledge distillation process serves as the core mechanism within the FedSDAF framework, designed to facilitate effective knowledge sharing among various clients. This process enables the model to transfer and integrate knowledge across different source domains, thereby achieving a comprehensive utilization of knowledge from multiple sources. Specifically, the bidirectional knowledge distillation involves computing two Kullback-Leibler (KL) divergences using the following formulas:
$L_{di}^{KL}=\text{KL}(z_{di}(x)\parallel\hat{z}_{dw}(x))$ (3)
$L_{dw}^{KL}=\text{KL}(\hat{z}_{dw}(x)\parallel z_{di}(x))$ (4)
Here, KL denotes the Kullback-Leibler divergence, while $z_{di}$ and $\hat{z}_{dw}$ represent the prediction logits from the model injected with $A_{di}$ and $\hat{A}_{dw}$, respectively. Through this setup, bidirectional knowledge distillation allows clients to exchange knowledge, facilitating learning and adaptation across different source domains. By sharing prediction logits, the model is able to acquire additional domain information from the server pertaining to other domains. This interaction of information is crucial in compensating for knowledge gaps arising from data isolation.
During the actual training process, we employ a distillation loss in conjunction with Cross-Entropy Loss ($L^{CE}$), introducing a weighting factor $\alpha$ to maintain a balance between the two losses. Figure 3 illustrates this process of bi-directional distillation:
@@ III-C2 Bidirectional Knowledge Distillation (BKD)
III-C3 Optimization process
The optimization process unfolds across the server and clients. On the server, the global domain-invariant adapter $A_{di}^{0}$ is initialized. In each communication round $r$, the server distributes the current global adapter $A_{di}^{r}$ to a subset of clients. Each participating client $k$ receives $A_{di}^{r}$ and performs local training on its dataset $\mathcal{D}_{k}$. During local training, the client updates its local copy of the domain-invariant adapter $\hat{A}_{di}^{k,r}$, and its private domain-aware adapter $A_{dw}^{k}$, using the bidirectional knowledge distillation loss described in Equation (5). After local training, each client sends its updated adapter $\hat{A}_{di}^{k,r}$ back to the server. The server then aggregates the received adapters to produce the next round’s global adapter $A_{di}^{r+1}$, using standard federated averaging:
@@ III-C3 Optimization process
where $n_{k}$ is the number of samples on client $k$ and $N=\sum_{k}n_{k}$ is the total number of samples across all clients. This iterative process allows the framework to build a robust global consensus within $A_{di}$ while leveraging the rich, localized knowledge preserved within each client’s $A_{dw}$.
The complete training process of FedSDAF unfolds iteratively across the server and clients, as formalized in Algorithm 1.
@@
1:Require: $K$: Number of clients; $\{D_{k}\}_{k=1}^{K}$: Client datasets; $R$: Communication rounds
2:Ensure: $\{A_{dw}^{k,R}\}_{k=1}^{K}$: Well-trained client models; $A_{di}^{R}$: Global domain-invariant adapter
@@
This iterative process allows the framework to build a robust global consensus within the domain-invariant adapter $A_{di}$ while continuously enriching it with the diverse, localized knowledge preserved within each client’s private domain-aware adapter $A_{dw}$.
IV Experimental Results
IV-A Datasets and Baselines
  1. 1.
  2. 2.
  3. 3.
TABLE II: Performance Comparison on PACS, VLCS, and OfficeHome Datasets (%)
Methods PACS VLCS OfficeHome
A C P S Avg. C L V S Avg. A C P R Avg.
Centralized learning based Domain Generalization Methods
CNN-based FedDG Methods (backbone: Resnet-50)
PEFT Methods (backbone: ViT-B/16)
TABLE III: Performance Comparison on DomainNet Datasets (%)
Methods DomainNet
C I P Q R S Avg.
FedCLIP $74.12$ $48.36$ $68.49$ $31.73$ $80.52$ $58.62$ $60.31$
PromptFL $76.53$ $51.72$ $70.86$ $34.21$ $81.68$ $68.37$ $63.90$
FedAPT $77.02$ $51.45$ $70.36$ $49.62$ $86.64$ $68.43$ $67.25$
FedPR $75.49$ $51.96$ $71.42$ $35.98$ $82.67$ $69.43$ $64.49$
FedMaPLe $78.61$ $65.23$ $71.89$ $43.46$ $86.32$ $72.46$ $69.67$
PLAN $79.51$ 66.42 $72.11$ $48.83$ 86.72 $72.69$ $71.05$
FedSDAF 82.59 $62.95$ 85.12 54.16 $84.99$ 74.14 73.99
IV-B Implementation Details
In our implementation of FedSDAF, we employed the ViT-B16 architecture as the backbone network. The federated learning process was configured to span 200 communication rounds, with each client performing 30 local epochs per round. To accommodate the diverse characteristics of the benchmark datasets, we utilized dataset-specific batch sizes: 128 for both the PACS and OfficeHome datasets, 64 for VLCS, and 1024 for the more extensive DomainNet. For optimization, we adopted the Scaffold optimizer with an initial learning rate of 0.001 to facilitate effective coordination of client model updates. To regulate the learning rate throughout the training process, we implemented PyTorch’s StepLR scheduler with a step size of 0.2 and a decay factor of 0.1, ensuring appropriate learning rate adjustments during model convergence.
IV-C Main Results
Tables II and III show the comparison of our proposed FedSDAF method with other methods on PACS, VLCS, OfficeHome, and DomainNet, respectively.
In our experimental evaluations, we achieved state-of-the-art (SOTA) results across all four mainstream datasets: PACS, VLCS, OfficeHome, and DomainNet. This accomplishment underscores the robust generalization capabilities of our proposed Federated Source Domain Awareness Framework (FedSDAF). Notably, our method demonstrated significant advancements on the Sketch domain of the PACS dataset, where previous approaches had struggled to deliver satisfactory outcomes. Specifically, our framework achieved an impressive accuracy of 96.56%, reflecting a notable improvement of 4.22% over prior methods.
IV-D Ablation Experiment
TABLE IV: Ablation Study for Different Components
BKD MHSA $A_{di}$ $A_{dw}$ PACS OfficeHome VLCS
98.52 87.63 91.38
$\times$ 97.46 85.81 86.19
$\times$ 97.13 85.42 87.20
$\times$ $\times$ 95.47 83.60 83.98
$\times$ $\times$ $\times$ 95.46 82.17 83.51
$\times$ $\times$ $\times$ $\times$ 93.29 78.70 82.69
To illustrate the significance of the various components utilized in the Federated Source Domain Awareness Framework (FedSDAF), we conducted an ablation study across three benchmark datasets. The results are presented in Table IV.
Optimal model performance was attained when all components were active, demonstrating the synergistic integration of features that collectively enhance the model’s generalization capacity. In contrast, disabling both the BKD and the MHSA led to a significant decline in performance, highlighting the critical roles these components play in facilitating information exchange and feature extraction. The observed drop in accuracy indicates that, without these mechanisms, the model struggles to maintain the understanding necessary for effective generalization. This finding underscores the necessity of collaborative learning strategies, such as BKD, in optimizing model performance within federated contexts characterized by data isolation.
IV-E Convergence Analysis
To validate the convergence of our FedSDAF framework, we compared it against the FedAvg (Adapter) baseline on the PACS dataset. As shown in Figure 4, FedSDAF converges significantly faster and more stably. Within the first 10 rounds, FedSDAF already achieves a 5.8–7.2% accuracy advantage. This rapid start is driven by the local domain-aware adapter’s MHSA mechanism, which accelerates the extraction of domain-specific features, whereas FedAvg’s domain-agnostic aggregation lags behind. By round 40, FedSDAF’s lead widens to a stable 9.4–13.8% accuracy margin. This superior stability is a direct result of our bidirectional knowledge distillation, which systematically fuses complementary knowledge across domains. This dynamic process overcomes the limitations of static aggregation strategies that are typically hindered by domain heterogeneity, ensuring a consistently faster and more robust convergence.
IV-F Hyperparameter Analysis
Figure 5: Effects of $\alpha$.
We evaluate the robustness of FedSDAF through systematic analysis of the loss balancing coefficient $\alpha$ (Eq. 5) across the entire PACS benchmark. As shown in Figure 5, FedSDAF consistently achieves accuracy exceeding 97% when $\alpha$ ranges from 0.3 to 5, with minimal variance ($\pm$0.8%). Notably, even with $\alpha=0.1$ (representing reduced distillation influence), the model maintains 96.2% accuracy, demonstrating the inherent stability of our dual-adapter architecture. Optimal performance is observed at $\alpha=1$, where domain-specific knowledge preservation and cross-client fusion reach an equilibrium. When $\alpha$ exceeds 5, a slight performance decline occurs (97.1% at $\alpha=10$), as the over-strengthened global distillation signal marginally inhibits local adaptation capabilities. These experimental results confirm FedSDAF’s robustness to hyperparameter selection, requiring only the straightforward choice of $\alpha=1$ to achieve peak performance without extensive tuning.
IV-G Communication Cost Analysis
The communication cost per round is dominated by two components: the upload/download of the lightweight domain-invariant adapter $\hat{A}_{di}$, and the exchange of logits for the BKD process. As shown in Figure 6, the size of the logits is negligible (0.003 MB). The total estimated communication cost for FedSDAF per client per round is approximately 2.270 MB. This is not only comparable to other PEFT-based methods like FedCLIP (2.010 MB) but is also an order of magnitude smaller than full-model aggregation methods like FedAvg (13.650 MB). This analysis demonstrates that FedSDAF remains highly communication-efficient, highlighting its practical potential for deployment in real-world federated learning scenarios.
IV-H Further Analysis
IV-H1 Fine-grained Ablation Study on Hard Domains